Jenelle has a piece of ribbon that measures 2 yards 8 inches in length she cuts the revenant before you go places what is the length in inches of each piece of ribbon
1
Expert's answer
2019-05-15T13:51:49-0400

2 yd = 6 ft = 72 in.

2 yd 8 in = 72+8 = 80 in.

80/2 = 40 in = 3 ft 4 in = 1 yd 4 in.